OpenStudy (anonymous):

How many silver (Ag) atoms are in a silver bar that is 20.0 g? A. 1.8 3 1023 Ag atoms B. 1.1 3 1023 Ag atoms C. 6.02 3 1023 Ag atoms D. 1.3 3 1027 Ag atoms

OpenStudy (zpupster):

\[\left( \frac{ 20g }{ 1} \right)\left( \frac{ 1 molAg }{107.8682g} \right)\left( \frac{ 6.023X10^{23} }{ mol } \right)\]

OpenStudy (anonymous):

I'm going with C, I'm not 100% on that though...

OpenStudy (ashley1nonly):

1.12*10^23

OpenStudy (anonymous):

B?

OpenStudy (zpupster):

yes